– 3(5 – 6) – 4(2 – 3) =
Found 2 solutions by rfer, nyc_function:
Answer by rfer(16322)

-3(5-6)-4(2-3)=
-3(-1)-4(-1)=
3+4=+7

Answer by nyc_function(2741)

– 3(5 – 6) – 4(2 – 3) =
-15 + 18 - 8 + 12 =
-23 + 30 = 7